OpenStudy (anonymous):

A proposed null hypothesis states that there is no difference between the mean weights of the population of two neighboring districts. The sample mean difference is 12 kilograms, and the standard deviation of the distribution of the difference between sample means is 9 kilograms. Which statement is true? A.The null hypothesis must be accepted if we choose the 68% confidence level. B.The null hypothesis must be accepted if we choose the 95% confidence level. C.The null hypothesis must be rejected if we choose the 99.7 % confidence level. D.The null hypothesis must be rejected if we

OpenStudy (anonymous):

choose the 100% confidence level. E.The null hypothesis must be rejected if we choose the 95% confidence level.

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

and its statistics..

OpenStudy (anonymous):

@shkrina

OpenStudy (anonymous):

@RANE

OpenStudy (uri):

Oh my I'm new to this subject,at this moment i can only be a student in this subject not a teacher but wair @hartnn and @ganeshie8 @amistre64 @whpalmer4 are pretty good,They pay help you.

OpenStudy (anonymous):

pay?

OpenStudy (uri):

may*

OpenStudy (anonymous):

oh alright

OpenStudy (rane):

i think its D bt not sure though

OpenStudy (anonymous):

is there a formula?

OpenStudy (rane):

no i didnt use any formula and neither do i think there is any for this

OpenStudy (anonymous):

so are you like 50% sure its that or is it just a guess

OpenStudy (rane):

70% sure

OpenStudy (anonymous):

alright, thank you(:

OpenStudy (rane):

yw:)

OpenStudy (anonymous):

@AccessDenied @mathslover

OpenStudy (anonymous):

?

OpenStudy (anonymous):

why did you tag people?

OpenStudy (anonymous):

@amistre64 because they may help u

OpenStudy (anonymous):

im fine now

OpenStudy (amistre64):

12/9 = 4/3 = 1.33 this tells us that 12 is 1.33 standard deviations from the mean

OpenStudy (amistre64):

the answer options suggest the empirical rule 68 is within 1 sd 95 is within 2 sd 99 is within 3 sd we reject the null if we are outside of our boundaries: 1.33 is greater then 1, so we reject it 68 Clevel 1.33 is less then 2, so we fail to reject it 95 Clevel 1.33 is less then 3, so we fail to reject it 99 Clevel

OpenStudy (anonymous):

your so good at stat lol

OpenStudy (amistre64):

youve given me alot of practice ;)

OpenStudy (anonymous):

help me wit my newest problem please (: